How to make 1M NaCl solution?

        1 mole NaCl = f. wt of NaCl = 58.5 g

        Take 58.5 g of NaCl and add water up to 1L-

H.W Q. 5.0g of Toluene(C7 H8 ). + 225 g of benzene (C6 H6 ).

             Density of solution is 0.876 g/mL.   Calculate Molarity

H.W Q. How many moles of HCl are in 45mL of 0.3 M HCl
Answer #1

Q1.

Note that the major component in the liquid solution is the solvent and the minor component is the solute.

Hence, in our mixture of toluene and benzene, benzene is the solvent and toluene is the solute.

Amount of toluene in the solution = 5.0 g

Molar mass of toluene ( C7H8) = 92.14 g/mol

Hence, number of moles of toluene in the solution is

5.09 n = mass molar mass 0.054 mol 92.14 g/mol

Molarity is defined as the number of moles of solute per litre of solution.

Mass of the solution = mass of toluene + mass of benzene = 5.0 g + 225 g = 230 g

Density of solution = 0.876 g/mL

Hence, the volume of the solution can be calculated as

mass density volum 30 g 0.876 g/mL =- volume >> volume = nome 0.876 g/mL 230 g 262.6 mL = 0.2626 L 202.

Now, the molarity can be calculated by dividing the number of moles of solute with the volume of solution in litres.

Hence,

n molarity = v= 0.054 mol 0.2626 L 0 2067 mol/L 0.207 M

Hence, the molarity of the solution is 0.207 M ( rounded to 3 significant figures).

Q2.

Concentration of the HCl solution = 0.3 M = 0.3 mol/L

Hence, there are 0.3 mol of HCl in 1000 mL of solution.

Hence, the number of moles of HCl in 45 mL can be calculated as follows:

0.3 mol x 45 mL = 0.0135 mol 1000 ml

Hence, there are 0.0135 mol of HCl in 45 mL of 0.3 M HCl solution.